# Understanding Iterative Combinatorial Auction Designs via Multi-Agent Reinforcement Learning

### v1

@@ 3.3.1. Monte Carlo Counterfactual Regret Minimization
Monte Carlo Counterfactual Regret Minimization (MCCFR) is a tabular MARL algorithm. It is effectively a generalization of regret matching to imperfect information extensive form games: at each information state, it tracks a regret for each action, playing actions with probabilities proportional to their regret. Specifically, we use a variant known as External Sampling MCCFR (Lanctot et al., 2009), which avoids traversing the entire game tree by randomly sampling chance outcomes and opponent actions. Initial experimentation revealed that two common performance tricks were helpful: we use the regret matching plus (Tammelin, 2014) algorithm instead of regret matching, and we use linear CFR (Brown and Sandholm, 2019a), which discounts regrets over time so that later iterations have larger impacts. Lastly, as is common for CFR, we analyze the average strategy across all iterations, rather than the last iterate.
We make one modification to this standard algorithm. To help steer MCCFR away from brittle equilibria, we train against trembling opponents. Specifically, when we sample an opponent’s action, we assume that they tremble 1% of the time, picking an action uniformly at random rather than using their trained policy. [^7] We do not make the training player tremble, as external sampling MCCFR already evaluates every one of their own actions rather than sampling just one. This trembling only occurs during training; there is no trembling in our final policies.
@@ 3.3.2. Proximal Policy Optimization
Our second algorithm is a deep reinforcement learning algorithm, Proximal Policy Optimization (PPO) (Schulman et al., 2017). PPO is an on-policy algorithm inspired by classic policy gradient methods. It represents its policy with a neural network; roughly, to train this policy, it collects a sample of information states, actions, and their eventual rewards, then updates its policy to increase the probability it places on actions leading to high rewards. PPO differs from a standard policy gradient method in that its objective function penalizes large changes in action probability, causing it to change its policy more slowly during training.
@@
In order to use PPO, we must choose a neural network architecture. We experiment with two. The first is a feed-forward neural network, accepting a flat vector describing all variables known in the information state as input. While this is a simple architecture, it is unlikely to be effective: such a network would have to learn how to process facts about every single bundle of licenses independently. To address this issue, we also test an equivariant architecture (Hartford et al., 2018), which is shown in Figure 1. In this architecture, each bundle has an associated set of features, and the network must perform the same computation on every bundle. Comparisons can be made across bundles using max-pooling. This enforces that the network must apply the same function to each bundle, giving it a strong inductive bias toward learning a single, simple function and applying it globally.
For both MCCFR and PPO, we convert all trained policies to pure strategies, placing probability 1 on the modal action. This conversion is computationally important, making it possible to compute NashConv by pruning all actions with probability 0.
@@ 3.4. Validating and Interpreting Policies
We compute several metrics about each of our trained policies. For each simulation, we record the auction revenue, welfare, length (in rounds), the number of unsold licenses, and number of lotteries in bid processing. The latter is relevant to the design decision being studied as it shines light on whether agents are engaging in strategies that trigger lotteries or carefully avoiding them. We validate the convergence of our modal policies by computing NashConv with a standard depth-first search algorithm.
We also compare our results to simulations of straightforward bidders. Concretely, we define a straightforward bidder as a bidder who, in each round, bids for the bundle of licenses that would give the highest utility at the current prices.  [^8] While we refer to this strategy as straightforward, we note that it is far from straightforward to define straightforward bidding in this auction. Any bid that involves dropping licenses might be rejected. This is especially important when part—but not enough—of a drop bid goes through and a bidder does not have enough activity to pick up what they intended. A bidder may therefore not want to place bids that could leave them exposed. Reasoning about bids in this manner requires a more complex demand forecasting model. This is a natural definition of straightforward bidding behavior as it is myopic, with no regard to price increases or activity constraints in future rounds. It is also analogous to existing definitions of straightforward bidding in simultaneous multi-round auctions (e.g., Milgrom, 2004), a related auction format.
@@ 4. Experimental Results
Our computational environment is described in Appendix A. Except where noted otherwise, we ran our MARL algorithms for one hour of walltime.
@@ 4.1.1. MCCFR: Trembling Opponents and Secondary Rewards
MCCFR has few hyperparameters that need to be tuned. Here, we evaluate our recommendations from Sections 2.2.2 and 2.2.4, where we advocated for breaking ties between rewards for actions, and for having opponents tremble to avoid brittle strategies. To test these techniques, we compared four variants of MCCFR, varying whether opponents trembled during training and whether the game included secondary rewards. We ran each variant with 10 random seeds on 10 different five-type games (comprising of 5 valuation samples and 2 bid processing rules), making for a total of 100 runs for each variant. For each run, we computed NashConv at six checkpoints, after 100, 200, 500, 1000, 2000, and 3000 iterations, respectively. Runs took between 10 and 60 minutes.
@@ 4.1.2. PPO: Architecture and Hyperparameter Optimization
Compared to CFR, PPO has many hyperparameters and its performance is highly sensitive to these hyperparameter settings. These hyperparameters control, among other elements, its neural network architecture, loss function, and optimizers. We tuned these hyperparameters using random search. Specifically, we sampled 60 random configurations of these hyperparameters and ran each with 3 random seeds on 6 different five-type games (comprising of 3 valuation samples and 2 bid processing rules), making for a total of 18 runs for each configuration. For each, we ran PPO for 30 minutes, computing NashConv once at the end. We provide details about the hyperparameters and our random sampling distributions in Appendix D.
The results are shown in Figure 3. The hyperparameter configurations varied substantially in their ability to find approximate equilibria, with many frequently producing policies with very high NashConv. However, a few configurations consistently converged to approximate equilibria; the best configuration always returned policies with a NashConv below 0.1. Notably, this configuration selected the AuctionNet architecture (as do 8 of the best 10 configurations). We used this single best PPO configuration in the remainder of our experiments.
@@ 4.2. Effects of Bid Processing Algorithms
We now turn to our main economic question: how does the auction’s bid processing algorithm affect bidders’ strategies and outcomes of the auction? To answer this question, we experimented with bidder valuations with 1, 2, 3, 5, and 7 types, generating 5 samples of each for a total of 25 value profiles. [^9] Bidders’ valuations were sampled independently for each of these games. This means that, for example, bidders’ values in the 1-type games were not necessarily equal to any values in games with more than 1 type. For each value profile, we created two games: one with each bid processing algorithm. These games varied in size, with the 1-type games having 10–100 information states, and the 7-type games having 500-700 information states. We ran MCCFR and PPO on each game with 10 random seeds, for a total of 1000 MARL algorithm runs. We discarded 24 runs that had a NashConv greater than 0.1, leaving 976 runs; 904 of these had a NashConv of 0, meaning that they converged to Nash equilibria.
@@
Our results are shown in Figure 4. This plot shows four metrics: the length of the auction, number of lotteries (i.e., the number of rounds where the bid processing algorithm can produce two or more different outcomes), revenue, and welfare. These metrics were averaged over ten thousand episodes to account for randomness in bidders’ assigned types and bid processing outcomes. Each point represents an (approximate) equilibrium found by one of the two MARL algorithms. We provide supplementary plots separating the equilibria found by each of the two MARL algorithms in Appendix E.
Our main finding is that the two bid processing rules lead to qualitatively different outcomes: when bidders have 3 or more types, equilibria under the drop-by-license rule can produce longer auctions with higher revenue and lower welfare than under the drop-by-player rule. These differences in auction outcomes are caused by differences in bidder behavior. Under drop-by-player, when both bidders attempt to switch their demand from the encumbered product to the unencumbered one, the bid processing algorithm can result in a lottery; when this happens, the auction immediately ends. Under drop-by-license, however, these lotteries are less desirable to bidders, as they risk dropping a single encumbered license which will not create enough activity to buy an unencumbered license. Bidders adjust their strategies accordingly, taking care to avoid simultaneously switching their demand, causing every single equilibrium under this rule to have no lotteries. In many games, this adjustment can lead to longer auctions, creating additional revenue as the prices increase.
@@
We also simulated the auctions under straightforward bidding. Straightforward simulation results for 7-type games are shown in Figure 5, with additional plots for the other games in Appendix E. Figure 5 shows that ignoring bidders’ incentives would lead to entirely different conclusions. Here, both rules lead to auctions of the same length, as bidders act the same way regardless of the bid processing algorithm. Furthermore, it is now the drop-by-player rule that leads to more revenue: under the drop-by-license rule, it is common for a license to be left unsold, losing revenue. The drop-by-player rule also produces higher welfare for a similar reason, as it is inefficient to leave licenses unsold. We note that straightforward bidding is indeed not an equilibrium in these auctions: on the 7-type games, NashConv ranges from 0.18 to 3.10.
@@ 4.3. Scaling to Larger Games
Encouraged by how well our algorithms performed on the games above, we created a set of larger games to test how well our MARL algorithms would scale. To do so, we added a third bidder and a fifth encumbered license. To keep the games tractable, we also increased the clock increment from 5% to 20%. We initialized the auction in a similar state as before, having each bidder bid for three encumbered licenses for two rounds. We generated 5 value profiles with 3 types per bidder, yielding a total of 10 games across the two bid processing algorithms. These games are substantially larger than the 2-player games: the smallest have around 8,000 information states, while the largest have over 40,000. We ran MCCFR for 8 hours on each game.
@@
Figure 6. (a) NashConv runtime distribution; (b) auction outcomes using MCCFR on 3-player games.
Validating trained policies on these larger games was substantially more difficult. At the end of each run, we attempted to compute NashConv for up to 1 hour. Within this time limit, we were only able to compute NashConv for 46 of the 100 runs; the distribution of runtimes, plotted in Figure 6a, shows that NashConv took at least 30 minutes on most runs. Nonetheless, when we were able to compute NashConv, it was low. The worst run’s NashConv was 0.03 and 36 runs had NashConvs of 0. Thus, it is possible for MCCFR to find equilibria of these larger games.
For completeness, Figure 6b shows the auction metrics for all 100 runs, including those for which we could not compute NashConv. It is interesting that the differences between the rules in this setting are quite unlike those in the 2-player games, with the drop-by-player rule leading to longer auctions, higher revenue, and higher welfare. One potential explanation is that under the drop-by-player rule, all three players can safely attempt to switch their demand from the encumbered product to the unencumbered product, allowing for further competition in the following round.
@@ 5. Conclusions and Discussion
Multiagent reinforcement learning algorithms are rapidly maturing, having recently reached several significant milestones. This paper argues that—if games are modelled appropriately and the right algorithmic choices are made—MARL can be used to analyze iterative combinatorial auctions that resist both pen-and-paper and computational Bayes–Nash analysis. We laid out a methodology for modelling an auction for use with MARL analysis, finding near-equilibria using MARL algorithms, and both validating and interpreting the results. We then offered a case study on bid processing in clock auctions, showing that MARL analysis could reach an economically meaningful conclusion in a setting that would not have admitted equilibrium analysis, and for which a simulation-based analysis would have yielded the wrong conclusion, and finding evidence that MARL analysis has the potential to scale to even larger settings.
We hope that we have persuaded the reader that MARL has something to offer auction analysis; however, our own work has clearly only scratched the surface. We thus conclude with an in-depth discussion of useful next steps: future clock auction design choices that would be fruitful to analyze; promising avenues for improving our methodology; and open theoretical questions.
@@ 5.1. Promising Auction Design Questions
A designer chooses the exponential increment by which prices of overdemanded products increase. Larger increments allow for quicker auctions, but reduce opportunities for price discovery as there is less feedback for bidders. Bidders react to these increments; in a consultation for the 3800 MHz Canadian auction, Telus “argued that increments [of 10%-20%] would create an accelerated cadence at the start of the auction that may prove contrary to the desired intention of promoting price discovery” (ISED, 2022). An auctioneer may even consider dynamic schemes, such as modifying the clock speed partway through the auction, or using a separate clock speed for each product, with steeper price increases for more overdemanded products.
@@ 5.2. Methodological Improvements
Our methodology recommended identifying multiple equilibria by running MARL algorithms with different random seeds. This approach is simple to implement, but may not be able to find all equilibria of a game; furthermore, inductive biases in the algorithm almost certainly make it more likely to find some equilibria than others. While this may be seen as a good thing if one believes that equilibria that learning dynamics struggles to find are less plausible, it may also be seen as a blind spot. An alternate approach that could be worth investigating is termed quality diversity in single-agent RL (e.g., Pugh et al., 2016); it focuses on finding multiple distinct, good policies. Future work should develop and evaluate algorithms to increase the diversity of equilibria, expanding our ability to fully understand the equilibria of an auction.
@@
Our analysis in our case study focused mostly on outcomes. Particularly for bidder-focused “playbook” applications, it can be important to derive insights about what constitutes a good strategy. This requires that learned policies not only be measurable but be interpretable. Bertsimas and Paskov (2022) use CFR to train poker agents and then fit decision trees to the CFR strategies through supervised learning to produce interpretable poker strategies. Future work could similarly focus on distilling bidding behavior. A set of composable bidding strategy building blocks could be valuable both for understanding auctions and as heuristics for scale.

### v2

@@ 3.3.1. Monte Carlo Counterfactual Regret Minimization
Counterfactual Regret Minimization (CFR) is a tabular MARL algorithm. It is effectively a generalization of regret matching to imperfect information extensive form games: at each information state, it tracks a regret for each action, playing actions with probabilities proportional to their regret. Specifically, we use a variant known as External Sampling Monte-Carlo CFR (Lanctot et al., 2009), which avoids traversing the entire game tree by randomly sampling chance outcomes and opponent actions. Initial experimentation revealed that two common performance tricks were helpful: we use the regret matching plus (Tammelin, 2014) algorithm instead of regret matching, and we use linear CFR (Brown and Sandholm, 2019a), which discounts regrets over time so that later iterations have larger impacts. Lastly, as is common for CFR, we analyze the average strategy across all iterations, rather than the last iterate.
We make one modification to this standard algorithm. To help steer MCCFR away from brittle equilibria, we train against trembling opponents: when we sample an opponent’s action, we pick an action uniformly at random 1% of the time, and otherwise sample from their trained policy. This trembling only occurs during training; there is no trembling in our final policies.
@@ 3.3.2. Proximal Policy Optimization
Our second algorithm is a deep reinforcement learning algorithm, Proximal Policy Optimization (PPO) (Schulman et al., 2017). PPO is an on-policy algorithm inspired by classic policy gradient methods. It represents its policy with a neural network. Roughly, to train this policy, it collects a sample of information states, actions, and their eventual rewards, then updates its policy to increase the probability on actions leading to high rewards. PPO differs from a standard policy gradient method in that its objective function penalizes large changes in action probability, causing it to change its policy more slowly during training.
@@
In order to use PPO, we must choose a neural network architecture. We experiment with two. The first is a feed-forward neural network, accepting a flat vector describing all variables known in the information state as input. While this is a simple architecture, it is unlikely to be effective: such a network would have to learn how to process facts about every single bundle of licenses independently. To address this issue, we also test an equivariant architecture (Hartford et al., 2018), which is shown in Figure 1. In this architecture, each bundle has an associated set of features, and the network must perform the same computation on every bundle. Comparisons can be made across bundles using max-pooling. This enforces that the network must apply the same function to each bundle, giving it a strong inductive bias toward learning a single, simple function and applying it globally. We list the variables given to these neural networks in Appendix D.
For both MCCFR and PPO, we convert all trained policies to pure strategies, placing probability 1 on the modal action. This conversion is computationally important, making it possible to compute NashConv as described in Section 2.3.1.
@@ 3.4. Validating and Interpreting Policies
We compute several metrics about each of our trained policies. For each simulation, we record the revenue, welfare, length (in rounds), number of unsold licenses, and number of lotteries in bid processing. The latter is relevant to the design decision being studied, demonstrating whether agents are engaging in strategies that trigger lotteries or avoiding them. To validate convergence, we compute NashConv with a standard depth-first search algorithm.
We also compare our results to simulations of straightforward bidders: bidders who, in each round, bid for the bundle with the highest utility at the current prices. This is a natural definition of straightforward bidding behavior as it is myopic, with no regard to price increases or activity constraints in future rounds. [^7] In particular, this definition of straightforward bidding is ignorant of the fact that any bid to lower demand could be rejected in part or in full by the bid processing algorithm, leading to exposure. It would be interesting to extend this heuristic to incorporate a demand forecasting model, predicting whether a drop bid is likely to be processed. It is also analogous to existing definitions of straightforward bidding in simultaneous multi-round auctions (e.g., Milgrom, 2004), a related auction format.
@@ 4. Experimental Results
Our computational environment is described in Appendix A. Except where noted otherwise, we ran our MARL algorithms for one hour of walltime, computing NashConv once at the end of the run.
@@ 4.1.1. MCCFR: Trembling Opponents and Secondary Rewards
MCCFR has few hyperparameters that need to be tuned. Here, we evaluate our recommendations from Sections 2.2.2 and 2.2.4, where we advocated for breaking ties between rewards for actions, and for having opponents tremble to avoid brittle strategies. To test these techniques, we compared four variants of MCCFR, varying whether opponents trembled during training and whether the game included secondary rewards. We ran each variant with 10 random seeds on 10 different 5-type games (comprising of 5 valuation samples and 2 bid processing rules): a total of 100 runs for each variant. For each run, we computed NashConv at six checkpoints. Runs took between 10 and 60 minutes.
@@ 4.1.2. PPO: Architecture and Hyperparameter Optimization
Compared to CFR, PPO has many hyperparameters—controlling its neural network architecture, loss function, and optimizers—and its performance is highly sensitive to these hyperparameter settings. We tuned these hyperparameters using random search. Specifically, we sampled 60 random configurations of these hyperparameters and ran each with 3 random seeds on 6 different 5-type games (comprising of 3 valuation samples and 2 bid processing rules), making for a total of 18 runs for each configuration. For each, we ran PPO for 30 minutes, computing NashConv once at the end. We provide details about the hyperparameters and our random sampling distributions in Appendix E, including more extensive results.
The hyperparameter configurations varied substantially in their ability to find approximate equilibria, with many frequently producing policies with very high NashConv. However, a few configurations consistently converged to approximate equilibria; the best configuration always returned policies with a NashConv below 0.1. Notably, this configuration selected the AuctionNet architecture (as do 8 of the best 10 configurations). We used this single best PPO configuration in the remainder of our experiments.
@@ 4.2. Effects of Bid Processing Algorithms
We now turn to our main economic question: how does the auction’s bid processing algorithm affect bidders’ strategies and outcomes of the auction? To answer this question, we experimented with bidder valuations with 1, 2, 3, 5, and 7 types, generating 5 samples of each for a total of 25 value profiles. [^8] Bidders’ valuations were sampled independently for each of these games: e.g., bidders’ values in the 1-type games were not necessarily equal to any values in games with more than 1 type. For each value profile, we created one game for each bid processing algorithm. These games ranged in size from 10–100 information states for the 1-type games to 500–700 for the 7-type games. We ran MCCFR and PPO on each game with 10 random seeds, for a total of 1000 MARL algorithm runs. We discarded 24 runs with a NashConv greater than 0.1, leaving 976 runs; 904 of these had a NashConv of 0, meaning that they converged to Nash equilibria.
@@
Our results are shown in Figure 3. This plot shows four metrics: the length of the auction, number of lotteries (i.e., the number of rounds where the bid processing algorithm can produce two or more different outcomes), revenue, and welfare. These metrics were averaged over 10,000 episodes to account for randomness in bidders’ types and bid processing outcomes. Each point represents an (approximate) equilibrium found by one of the two MARL algorithms. We provide supplementary plots showing the equilibria found by each algorithm in Appendix F.
Our main finding is that the two bid processing rules lead to qualitatively different outcomes: when bidders have 3 or more types, equilibria under the drop-by-license rule can produce longer auctions with higher revenue and lower welfare than under the drop-by-player rule. These differences in auction outcomes are caused by differences in bidder behavior. Under drop-by-player, when both bidders attempt to switch their demand from the encumbered product to the unencumbered one, the bid processing algorithm results in a lottery; when this happens, the auction immediately ends. Under drop-by-license, however, these lotteries are less desirable to bidders, as they risk dropping a single encumbered license which will not create enough activity to buy an unencumbered license. Bidders adjust their strategies accordingly, taking care to avoid simultaneously switching their demand, causing every single equilibrium under this rule to have no lotteries. In many games, this adjustment can lead to longer auctions, creating additional revenue as the prices increase.
@@
We also simulated the auctions under straightforward bidding. Straightforward simulation results for 7-type games are shown in Figure 4, with additional plots for the other games in Appendix F. Figure 4 shows that ignoring bidders’ strategic incentives would lead to entirely different conclusions. Here, both rules lead to auctions of the same length, as bidders act the same way regardless of the bid processing algorithm. Furthermore, it is now the drop-by-player rule that leads to more revenue: under the drop-by-license rule, it is common for a license to be left unsold, losing revenue. The drop-by-player rule also produces higher welfare for a similar reason, as it is inefficient to leave licenses unsold. We note that straightforward bidding is indeed not an equilibrium in these auctions: on the 7-type games, NashConv ranges from 0.18 to 3.10.
@@ 4.3. Scaling to Larger Games
Encouraged by the results above, we created a set of larger games to test how well our MARL algorithms would scale. To do so, we added a third bidder and a fifth encumbered license. To keep the games tractable, we also increased the clock increment from 5% to 20%. We initialized the auction in a similar state as before, having each bidder bid for three encumbered licenses for two rounds. We generated 5 value profiles with 3 types per bidder, yielding a total of 10 games across the two bid processing algorithms. These games are substantially larger than the 2-player games, ranging from 8,000–40,000 information states. We ran MCCFR and PPO for 8 hours on each game.
@@
Figure 5. (a) NashConv runtime distribution; (b) auction outcomes, using MCCFR on 3-player games.
Validating trained policies on these larger games was substantially more difficult. At the end of each run, we attempted to compute NashConv for up to 1 hour. Within this time limit, we were only able to compute NashConv for 46 of the 100 MCCFR runs (Figure 5a). When we were able to compute NashConv, we found that MCCFR was consistently able to find (approximate) Nash equilibria: 36 runs had a NashConv of 0, and the highest computed NashConv was 0.03. Figure 5b shows the auction metrics for all 100 MCCFR runs, including those for which we could not compute NashConv. Notably, the differences between the rules in this setting are quite unlike those in the 2-player games, with the drop-by-player rule leading to longer auctions, higher revenue, and higher welfare. One potential explanation is that under the drop-by-player rule, all three players can safely attempt to switch their demand from the encumbered product to the unencumbered product, allowing for further competition in the following round.
We found that PPO performed far worse than MCCFR. We were able to compute NashConv for 37 of 50 PPO runs, finding that it was unable to reliably find approximate equilibria: almost all runs had a NashConv greater than 3, and the majority had a NashConv greater than 10. Thus, it appears nontrivial to scale PPO to larger games; we suspect that its hyperparameters would need to be retuned for this larger environment in order to perform well.
@@ 5. Conclusions and Discussion
Multiagent reinforcement learning algorithms are rapidly maturing, having recently reached several significant milestones. This paper argues that—with the right modeling and algorithmic choices—MARL can be used to analyze iterative combinatorial auctions that resist both pen-and-paper and computational Bayes–Nash analysis. We laid out a methodology for modelling an auction for use with MARL analysis, finding near-equilibria using MARL algorithms, and both validating and interpreting the results. We then offered a case study on bid processing in clock auctions, showing that MARL analysis could reach an economically meaningful conclusion in a setting that does not admit analysis with existing approaches, and for which a simulation-based analysis would have yielded the wrong conclusion, with potential to scale to even larger settings.
We hope that we have persuaded the reader that MARL has something to offer auction analysis; however, our own work has clearly only scratched the surface. We thus conclude with a discussion of useful next steps: future clock auction design choices that would be fruitful to analyze; promising avenues for improving our methodology; and open theoretical questions.
@@ 5.1. Promising Auction Design Questions
A designer chooses how quickly the prices of overdemanded products increase. Larger increments allow for shorter auctions, but give less feedback to bidders, reducing opportunities for price discovery. Bidders react to these increments: in a consultation for the 3800 MHz Canadian auction, Telus “argued that increments [of 10%-20%] would create an accelerated cadence at the start of the auction that may prove contrary to the desired intention of promoting price discovery” (ISED, 2022). An auctioneer may even consider dynamic schemes, modifying the clock speed throughout the auction, or using steeper price increases for more overdemanded products.
@@ 5.2. Methodological Improvements
5.2.1. Scaling to Larger Games.
A limitation of our experiments is their scale: even our larger games (Section 4.3) only had tens of thousands of information states, relatively small by reinforcement learning’s standards. However, scaling to larger games presents many challenges. As the game tree grows large, tabular methods require more space to store their policies; it becomes infeasible complete even a single iteration of a tree-traversal method; and computing NashConv exactly becomes unrealistic, complicating both analysis and hyperparameter tuning. This final point is especially impactful for deep learning methods, which typically need to have their hyperparameters re-tuned for different environments. Future work should study the extent to which these challenges can be overcome, scaling MARL algorithms to larger and more complex auctions.
@@
We found multiple equilibria by running MARL algorithms with different random seeds. This approach is easy to implement, but may not be able to find all equilibria of a game. Furthermore, inductive biases in the algorithm make it more likely to find some equilibria than others. While one might believe that equilibria which are hard to find are less plausible, this may also be seen as a blind spot. Future work should develop and evaluate algorithms to increase the diversity of equilibria, expanding our ability to fully characterize the equilibria of an auction. For example, it may be fruitful to adapt ideas from quality diversity in single-agent RL (e.g., Pugh et al., 2016), which aims to find multiple distinct, good policies.
@@
While our case study focused mostly on auction outcomes, it can also be important to derive insights about what constitutes a good strategy. This requires that learned policies be not only measurable, but interpretable. Bertsimas and Paskov (2022) fit decision trees to trained CFR strategies to produce interpretable poker strategies. Future work could similarly focus on distilling bidding behavior. A set of composable building blocks of bidding strategies could be valuable both for understanding auctions and to inform how to shrink bidders’ action space in larger-scale experiments.